Quadratic Function Solve: X squared + x - 3 = 0
Answer by alka(15)

x^2 + x - 3 = 0
a= 1 , b = 1 , c = -3
D = b^2 - 4ac
= 13
x = (-b+sqrt(D))/2a , x = (-b-sqrt(D))/2a
=(-1+sqrt(13))/2 , x = (-1-sqrt(13))/2